OpenStudy (d9414):

A​ store's selling price for a TV includes a​ 135% mark up over what the store paid for the TV. The store decides to discount their selling price by​ 40%. You have a coupon for​ 15% off the discount price. Find the percentage change in the​ store's cost for the TV.

OpenStudy (retireed):

x = $100 the store's cost for the TV x*1.35 + x = selling price $235 235 * 0.4 = discount off the selling price 94.0 235 - 94 = the on-sale price $141 141 * 0.15 = on-sale price coupon discount 21.15 141 - 21.15 = purchase price $119.85 Find the percentage change in the​ store's cost for the TV. I get 19.85% Does this makes sense, because I don't know if my methodology is correct? I will need to recheck my math with a calculator, if it doesn't seem right to you.

OpenStudy (d9414):

give me one second, I'm working out the problem myself the way you did to make sure i understand it

OpenStudy (d9414):

okay in the beginning I thought it was 100% plus 135%

OpenStudy (d9414):

@re

OpenStudy (retireed):

okay

OpenStudy (retireed):

That is roughly what I did. I just used a random TV price and used $100 as the store cost.

OpenStudy (d9414):

and how did you get to 19.85 after 119.85?

OpenStudy (retireed):

The original store cost was $100 and the price I bought it at is $119.85. This is the questionable part so the percentage change in the stores cost $119.85/$100 = 1.1985 - 1 = 0.1985 times 100 to make a percentage 19.85%

OpenStudy (d9414):

ohhh okay i see. thank you so much. can I post another problem I need help with?

OpenStudy (retireed):

No this one hurt my brain too much. You can post it and answer as much as you can and I will try to help you if I can follow it. The best way to work on this site is to post a question and show as much of your work as possible. We are here to guide you. On this problem I fail you because I didn't guide you I feel I did the whole thing. You would be better off with another mentor.